OpenStudy (anonymous):

Help Please! Which is the best equation to use to solve this problem? The perimeter of a rectangle is 68 ft. The ratio of the length to the width is 9 : 8. What are the dimensions of the rectangle? A. 9x+8x=68 B. 2x+2x=68 C. 9x+9x+8x+8x=68 D. 9x --- = 68 8x

OpenStudy (anonymous):

2 * L + 2 * W = 68 2*L = 2* 9x = 9x + 9x 2*W = 2*8x = 8x + 8x Since the perimeter is 68 9x + 9x + 8x + 8x = 68

OpenStudy (anonymous):

C

OpenStudy (anonymous):

you good?

OpenStudy (anonymous):

@chaser71 ok thanks :) can i get your help i am doing a 25 question test and need help with four more

OpenStudy (anonymous):

goo ahead! :D